what is sum of square of the roots of eq. 2x^2+3x+56

Grade:upto college level

1 Answers

Latika Leekha
9 years ago
Please note we need to have an equality in this to call it an equation.
The equation should have been 2x2+ 3x + 56 = 0.
We know that the sum of roots is given by -b/a and the product of roots is c/a.
here, a = 2, b = 3, c = 56.
Now, sum of roots = -b/a = -3/2
Product of roots = c/a = 56/2 = 28.
Let us assume the roots to be A and B.
Then A + B = -3/2 and AB = 28.
We are required to find A2 + B2.
A2 + B2 = (A+B)2 – 2AB
= (-3/2)2 – 2. 28.
= 9/4 – 56
= -215/4.
Thanks & Regards
Latika Leekha
askIITians Faculty